# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 5.2. Unipotent invariance in quotients
Let $\tilde{H}_{u}:=\text{zcl}(H_{u})$ with $H_{u}$ being the subgroup generated by all one-parameter unipotent subgroups contained in $H$. Since $H\trianglelefteq N$, it follows that $H_{u}\trianglelefteq N$, and by Zariski density we get that $\tilde{H}_{u}\trianglelefteq N$. Now, $H_{u}$ is of finite index in $\tilde{H}_{u}$, therefore, since $\psi$ is unbounded in $N/H_{u}$ it follows that $\psi$ is unbounded in $N/\tilde{H}_{u}$.
By [Springer_book, Theorem 5.5.3] and since $\tilde{H}_{u}$ has no non-trivial real character, there is a finite dimensional rational representation $\rho:\text{SL}(n,\mathbb{R})\to\text{GL}(V)$ with an $v\in V$ such that
$\tilde{H}_{u}=\{g\in\text{SL}(n,\mathbb{R}):\rho(g)v=v\}.$
Now let
$W:=\{x\in V:\rho(h)x=x,\forall h\in\tilde{H}_{u}\}.$
We have that $v\in W$ so that $W$ is not the trivial subspace. Because $\tilde{H}_{u}$ is a normal subgroup of $N$, we get that the action of $N$ keeps $W$ invariant. Now recall that one-parameter unipotent subgroups generate $H_{\psi}$. Therefore, $\det(\rho(H_{\psi})|_{W})=\{1\}$. In particular, $\det(\rho(\psi)|_{W})=\{1\}$.
Since $\psi$ is not contracting, we have in particular that $\lim_{t\to\infty}\rho(\psi(t))w\neq 0,~{}\forall w\in W\smallsetminus\{0\}.$ Thus, by Lemma 5.9, there is an $r<1$ such that
$\lim_{t\to\infty}\rho(\psi(h_{s,r}(t))\psi(t)^{-1})|_{W}=:\bar{p}(s),s\in\mathbb{R},$
where $\overline{p}(s)$ is a one-parameter unipotent subgroup. It is a general fact if $f:G_{1}\to G_{2}$ is a homomorphism of algebraic groups, and $U_{2}\leq G_{2}$ is a unipotent one-parameter subgroup, then there is a unipotent one-parameter subgroup $U_{1}\leq G_{1}$ such that $f(U_{1})=U_{2}$. This follows by Jordan decomposition. Thus, there is a unipotent one-parameter subgroup $U\leq N$ such that $\rho(U)|_{W}=\{\bar{p}(s):s\in\mathbb{R}\}$. Now consider the diagram:
[…]
Since images of algebraic groups under algebraic homomorhpisms are closed (see e.g. [Springer_book, Proposition 2.2.5]), we conclude that $\tilde{\tau}$ is an injective proper map. Since $H_{u}$ is of finite index in $\tilde{H}_{u}$, the map $\tau$ is a covering map with finite fibers, and it follows that $\tilde{\tau}\circ\tau$ is a proper map with finite fibers. Then, we obtain for each $s\in\mathbb{R}$,
(5.16) $\lim_{t\to\infty}\vartheta(\psi(h_{s,r}(t))\psi(t)^{-1})=\lim_{t\to\infty}\vartheta(\psi(h_{s,r}(t)))\vartheta(\psi(t))^{-1}=:\tilde{p}(s).$
Now we show that $\tilde{p}(s),s\in\mathbb{R}$ is a one-parameter group, where $s\mapsto\tilde{p}(s)$ is an homomorphism from the additive group of real numbers. This argument appears in [Poulios_thes] and given here for completeness. Note that for $r<1$ and for all $s,l\in\mathbb{R}$ that
$h_{s,r}\circ h_{l,r}=h_{s+l,r}.$
Denote $\xi(t):=\vartheta(\psi(t))$. Now if (5.14) holds for $r<1$, then for $s,l\in\mathbb{R}$ we have
$\tilde{p}(s)\tilde{p}(l)$ $=(\lim_{t\to\infty}\xi(h_{s,r}(t))\xi(t)^{-1})(\lim_{t\to\infty}\xi(h_{l,r}(t))\xi(t)^{-1})$
$=(\lim_{t\to\infty}\xi(h_{s,r}(h_{l,r}(t)))\xi(h_{l,r}(t))^{-1})(\lim_{t\to\infty}\xi(h_{l,r}(t))\xi(t)^{-1})$
$=\lim_{t\to\infty}\xi(h_{s,r}(h_{l,r}(t)))\xi(h_{l,r}(t))^{-1}\xi(h_{l,r}(t))\xi(t)^{-1}$
$=\lim_{t\to\infty}\xi(h_{s+l,r}(t))\xi(t)^{-1}=\tilde{p}(s+l).$
Since the above diagram commutes, we get that $\vartheta(U)=\{\tilde{p}(s):s\in\mathbb{R}\}$ is a non-trivial subgroup of $N/H_{u}$. Finally, consider the natural quotient map $\eta:N/H_{u}\to N/H$, and note that $q=\eta\circ\vartheta$. Let $p(s):=\eta(\tilde{p}(s))$. Then, due to (5.16), we get by continuity that for all $s\in\mathbb{R}$,
@@
Finally, since $U\not\subset H_{u}$ and since $U$ is unipotent, it follows that $q(U)\leq N/H$ is not trivial, and $q(U)=\eta\circ\vartheta(U)=\{p(s):s\in\mathbb{R}\}.$
5.3. Inclusion of the hull in a Lie group
Here, we will prove the following.
@@ 5.3. Inclusion of the hull in a Lie group
Let $G\leq\text{SL}(n,\mathbb{R})$ be a connected unimodular closed subgroup, and let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve, definable in a polynomially bounded o-minimal structure. Then, $H_{\varphi}\leq G_{u}$ where $H_{\varphi}$ is the hull of $\varphi$ and $G_{u}\leq G$ is the closed subgroup generated by the one-parameter unipotent subgroups contained in $G$. In particular, both the hull and the range of the correcting curve of $\varphi$ are in $G$.
@@
$\text{Trace}(\text{ad}X)=\text{Trace}(\text{ad}X|_{\mathfrak{g}_{u}}).$
@@
$\text{Trace}(T|_{V})=\text{Trace}(T|_{W})+\text{Trace}(T_{V/W}),$
@@
$\text{Trace}(\text{ad}X|_{\mathfrak{g}})=\text{Trace}(\text{ad}X|_{[\mathfrak{g},\mathfrak{g}]}).$
Let $\tilde{\mathfrak{g}}$ denote the Lie algebra of the Zariski closure of the Lie group associated with $\mathfrak{g}$ in $\text{GL}(n,\mathbb{R})$. By [Chevalley-LieII, Chapter II, Theorem 13], $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=[\mathfrak{g},\mathfrak{g}]$. Thus,
$\text{Trace}(\text{ad}X|_{\mathfrak{g}})=\text{Trace}(\text{ad}X|_{[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]}).$
Let $\mathfrak{u}$ denote the radical of the $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$. Then $\mathfrak{u}$ consists of nilpotent matrices; we can derive this from [RA72, 2.5 A structure theorem]. We also note that $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$ and $\mathfrak{u}$ are ideals in $\tilde{\mathfrak{g}}$. Since $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]/\mathfrak{u}$ is semisimple, the trace of the derivation on it corresponding to $\text{ad}X$ is zero. Therefore,
$\text{Trace}(\text{ad}X|_{[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]})=\text{Trace}(\text{ad}X|_{\mathfrak{u}}).$
Now let $\mathfrak{v}$ denote the radical of $\mathfrak{g}_{u}$. Since $\mathfrak{g}_{u}$ is a Lie subalgebra generated by nilpotent matrices, $\mathfrak{v}$ consists of nilpotent matrices. Also, $\mathfrak{g}_{u}$, and hence $\mathfrak{v}$ are ideals of $\mathfrak{g}$. As above, since $\mathfrak{g}_{u}/\mathfrak{v}$ is semisimple, we conclude that
$\text{Trace}(\text{ad}X|_{\mathfrak{g}_{u}})=\text{Trace}(\text{ad}X|_{\mathfrak{v}}).$
It remains to prove:
$\text{Trace}(\text{ad}X|_{\mathfrak{v}})=\text{Trace}(\text{ad}X|_{\mathfrak{u}}).$
Now $\mathfrak{u}\subset[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=[\mathfrak{g},\mathfrak{g}]\subset\mathfrak{g}$. Hence, $\mathfrak{u}\subset\mathfrak{g}_{u}$, and so $\mathfrak{u}\subset\mathfrak{v}$. Moreover, $\mathfrak{v}\cap[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$ is an ideal in $\tilde{\mathfrak{g}}$, so it is contained in $\mathfrak{u}$. Thus,
$\mathfrak{v}\cap[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=\mathfrak{u}.$
Therefore $[X,\mathfrak{v}]\subset\mathfrak{v}\cap[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=\mathfrak{u}$. Therefore, the action of $\text{ad}X$ on $\mathfrak{v}/\mathfrak{u}$ is zero. So,
$\text{Trace}(\text{ad}X|_{\mathfrak{v}})=\text{Trace}(\text{ad}X|_{\mathfrak{u}}).$
Corollary 5.15.
Let $G$ be a closed connected Lie subgroup of $\text{GL}(n,\mathbb{R})$. Suppose that $G$ is unimodular. Let $G_{u}$ be the closed subgroup generated by unipotent one-parameter subgroups contained in $G$ and let $\mathfrak{g}_{u}$ denote the Lie algebra of $G_{u}$. Then for any $g\in G$,
$\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=1.$
Proof.
Suppose that $g=\exp(X)$. Then,
$\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=\exp(\text{Trace}(\text{ad}X|_{\mathfrak{g}_{u}})).$
Then, Lemma 5.14 shows that $\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=\det(\text{Ad}(g))$. Since $G$ is connected and unimodular, we have that $\det(\text{Ad}(g))=1$. Then, $\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=1$ for all $g$ in some identity neighborhood in $G$. Since the identity neighborhood generates $G$, the corollary follows. ∎
@@
Let $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve, definable in a polynomially bounded o-minimal structure. Let $\beta$ be a correcting curve for $\varphi$, and denote $\psi(t):=\beta(t)\varphi(t)$.
Consider the exterior product of the Adjoint representation of $\text{SL}(n,\mathbb{R})$ on $\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{sl}(n,\mathbb{R})$. Let $p\in\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{g}_{u}\setminus\{0\}$, and let $N$ be the stabilizer of $p$ in $\text{SL}(n,\mathbb{R})$. Note that $N$ is Zariski closed, $N$ is observable, and $G_{u}\trianglelefteq N$. By Corollary 5.15, $G\subset N$. Since $N$ is observable, and $\varphi\subset N$, we conclude that $H_{\varphi}\subseteq N$ by the minimality of the hull, Theorem 2.4,(1). Now, by Remark 2.6, we have that $\psi\subset H_{\psi}=H_{\varphi}$. Assume for contradiction that $\psi$ is not bounded modulo $G_{u}$. Then, by Proposition 5.12, there exists a one-parameter unipotent subgroup of $\{u(s)\}_{s\in\mathbb{R}}\leq N$ such that
$q(u(s))=\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1},s\in\mathbb{R}.$
Also $q\circ u$ is a nontrivial unipotent one-parameter subgroup of $N/G_{u}$. Let $\beta_{\infty}=\lim_{t\to\infty}\beta(t)$. Then, for all $s\in\mathbb{R}$,
@@
and since $\varphi\subset G\subset N$, we get that $\{q(\beta_{\infty}^{-1}u(s)\beta_{\infty}):s\in\mathbb{R}\}$ is a nontrivial one-parameter subgroup of $G/G_{u}$. It then follows that $\{\beta_{\infty}^{-1}u(s)\beta_{\infty}:s\in\mathbb{R}\}$ is a nontrivial one-parameter unipotent subgroup contained in $G$ not contained in $G_{u}$, a contradiction. Finally, since $\varphi$ is bounded modulu $G_{u}$, and since $\text{zcl}(G_{u})$ is observable, we get that $H_{\varphi}\subseteq G_{u}$ by the minimality of the hull, Theorem 2.4,(1). ∎
6. Linearization
We begin with preliminaries needed for the linearization technique. Suppose that $\Gamma\leq\text{SL}(n,\mathbb{R})$ is a discrete subgroup, and let $\mu$ be a probability measure on $X:=\text{SL}(n,\mathbb{R})/\Gamma$. Suppose that there exists a subgroup $W\leq G$ of positive dimension which is generated by one-parameter unipotent subgroups contained in $W$ such that $\mu$ is invariant under the $W$-left translates. By the celebrated Ratner’s theorem (see [Ratner91a]) every $W$-ergodic component of $\mu$ is homogeneous.

### v2

@@ 5.2. Unipotent invariance in quotients
Let $\tilde{H}_{u}:=\text{zcl}(H_{u})$, where $H_{u}$ is the subgroup generated by all one-parameter unipotent subgroups contained in $H$. Since $H\trianglelefteq N$, it follows that $H_{u}\trianglelefteq N$, and by Zariski density $\tilde{H}_{u}\trianglelefteq N$. Now, $H_{u}$ is of finite index in $\tilde{H}_{u}$ (see Remark 2.7), therefore, since the image of $\psi$ is unbounded in $N/H_{u}$ it follows that the image of $\psi$ is unbounded in $N/\tilde{H}_{u}$.
We realize $N/\tilde{H}_{u}$ as a real algebraic subgroup of $\text{SL}(m,\mathbb{R})$ for some $m$. For each $t\in[0,\infty)$, let $\bar{\psi}(t)$ denote the image of $\psi(t)$ in $\text{SL}(m,\mathbb{R})$. We claim that $\bar{\psi}$ is non-contracting for $\text{SL}(m,\mathbb{R})$. Indeed, because any finite-dimensional linear representation of $\text{SL}(m,\mathbb{R})$ is an algebraic representation of $\tilde{H}_{\psi}\subseteq N$, and $\psi$ is non-contracting for $\tilde{H}_{\psi}$, where $\tilde{H}_{\psi}$ is the Zariski closure of the hull of $\psi$, see Theorem 2.5, (3).
Therefore by Lemma 5.9, there exists $r<1$ such that
(5.13) $\lim_{t\to\infty}\bar{\psi}(h_{s,r}(t))\bar{\psi}(t)^{-1}=:\bar{p}(s),\,\forall s\in\mathbb{R},$
where $s\mapsto\overline{p}(s)$ is a nontrivial one-parameter unipotent subgroup of $\text{SL}(m,\mathbb{R})$ contained in $N/\tilde{H}_{u}$. Let $\vartheta:N\to N/H_{u}$ denote the quotient homomorphism. Since the quotient map $N/H_{u}\to N/\tilde{H}_{u}$ is a homomorphism with a finite kernel, for each $s\in\mathbb{R}$, the following limit exists:
(5.14) $\tilde{p}(s):=\lim_{t\to\infty}\vartheta(\psi(h_{s,r}(t)))\vartheta(\psi(t))^{-1}.$
In view of Jordan decomposition, there exists a unipotent one-parameter subgroup $\{u(s):s\in\mathbb{R}\}\subseteq N$ such that for all $s\in\mathbb{R}$, the image of $u(s)$ in $N/\tilde{H}_{u}$ is $\bar{p}(s)$, and hence $\tilde{p}(s)=\vartheta(u(s))$. Since $\{\bar{p}(s):s\in\mathbb{R}\}$ is a unipotent group, we obtain that $\{u(s):s\in\mathbb{R}\}$ is a nontrivial unipotent one-parameter subgroup of $N$ not contained in $H_{u}$, and hence it is not contained in $H$. Thus, $\{q(u(s)):s\in\mathbb{R}\}$ is a nontrivial subgroup of $N/H$.
Denote $\eta:N/H_{u}\to N/H$ the quotient homomorphism, and note that $\eta\circ\vartheta=q$. Then, for all $s\in\mathbb{R}$, $p(s):=\eta(\tilde{p}(s))=q(u(s))$, and by (5.14),
@@
5.3. Inclusion of the hull in a Lie subgroup
@@
Let $G\subseteq\text{SL}(n,\mathbb{R})$ be a connected unimodular closed subgroup, and let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting, continuous curve such that $\varphi([0,\infty))\subseteq G$. Then, $H_{\varphi}\subseteq G_{u}$, and in particular, any correcting curve of $\varphi$ takes values in $G_{u}$.
The result does not follow from Theorem 2.5, because we do not assume that $G$ is an observable subgroup of $\text{SL}(n,\mathbb{R})$.
If $G$ is not algebraic, the hull $H_{\varphi}$ may be contained in the Zariski closure of $G$, but not in $G$. The unimodularity of $G$ ensures that $H_{\varphi}\subseteq G$.
The unimodularity assumption is essential even in the algebraic subgroup case. For example, in Remark 5.11, the curve is non-contracting in $\text{SL}(2,\mathbb{R})$ and its image is contained in the non-unimodular Borel subgroup $B$ of $\text{SL}(2,\mathbb{R})$. But its hull is not contained in $B$.
The following is an immediate corollary of Proposition 5.13.
Corollary 5.14.
Let $G_{1}\subseteq G_{2}$ be real points of real algebraic groups. Let $\varphi:[0,\infty)\to G_{2}$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting, continuous curve. Suppose that $\varphi([0,\infty))\subseteq G_{1}$ and $G_{1}$ is unimodular. Then $\varphi:[0,\infty)\to G_{1}$ is non-contracting.
@@
$\text{tr}(\text{ad}X)=\text{tr}(\text{ad}X|_{\mathfrak{g}_{u}}),$
where $\text{tr}(A)$ denotes the trace of $A$.
So, for the connected Lie subgroup $G$ corresponding to $\mathfrak{g}$, we have
(5.15) $\det(\text{Ad}(g)|_{\mathfrak{g}_{u}})=\det(\text{Ad}(g)),\ \forall g\in G.$
@@
(5.16) $\text{tr}(T|_{V})=\text{tr}(T|_{W})+\text{tr}(T_{V/W}),$
@@
(5.17) $\text{tr}(\text{ad}X|_{\mathfrak{g}})=\text{tr}(\text{ad}X|_{[\mathfrak{g},\mathfrak{g}]}).$
Let $\tilde{\mathfrak{g}}$ denote the Lie algebra of the Zariski closure of the Lie group associated with $\mathfrak{g}$ in $\text{GL}(n,\mathbb{R})$. By [Chevalley-LieII, Chapter II, Theorem 13],
$[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]=[\mathfrak{g},\mathfrak{g}]\subseteq\mathfrak{g}.$
Let $\mathfrak{u}$ denote the radical of the $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$. Then $\mathfrak{u}$ consists of nilpotent matrices; we can derive this from [RA72, (2.5) A structure theorem]. So, $\mathfrak{u}\subseteq\mathfrak{g}_{u}$.
Since $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]$ and $\mathfrak{u}$ are ideals in $\tilde{\mathfrak{g}}$ and $[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]/\mathfrak{u}$ is semisimple, the trace of the derivation on it corresponding to $\text{ad}X$ is 00. Therefore,
(5.18) $\text{tr}(\text{ad}X|_{[\tilde{\mathfrak{g}},\tilde{\mathfrak{g}}]})=\text{tr}(\text{ad}X|_{\mathfrak{u}}).$
Let $\mathfrak{v}$ denote the radical of $\mathfrak{g}_{u}$. Then, $\mathfrak{u}\subseteq\mathfrak{v}$. Since $\mathfrak{g}_{u}$ is a Lie subalgebra generated by nilpotent matrices, $\mathfrak{v}$ consists of nilpotent matrices. Also, $\mathfrak{g}_{u}$, and hence $\mathfrak{v}$ are ideals of $\mathfrak{g}$. So, by definition of $\tilde{\mathfrak{g}}$, $\mathfrak{v}$ is an ideal of $\tilde{\mathfrak{g}}$. So, $\mathfrak{v}\subseteq\mathfrak{u}$. Hence $\mathfrak{v}=\mathfrak{u}$.
As above, since $\mathfrak{g}_{u}/\mathfrak{v}$ is semisimple, we conclude that
(5.19) $\text{tr}(\text{ad}X|_{\mathfrak{g}_{u}})=\text{tr}(\text{ad}X|_{\mathfrak{v}}).$
Now, (5.16) follows from (5.17)-(5.19). ∎
@@
Consider the exterior product of the Adjoint representation of $\text{SL}(n,\mathbb{R})$ on $\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{sl}(n,\mathbb{R})$, and fix $p\in\wedge^{\dim\mathfrak{g}_{u}}\mathfrak{g}_{u}\setminus\{0\}$. Let
$N$ :=$:=\{g\in\text{SL}(n,\mathbb{R}):g\cdot p=p\}$
=$=\{g\in\text{SL}(n,\mathbb{R}):\text{Ad}(g)\mathfrak{g}_{u}=\mathfrak{g}_{u},\,\det(\text{Ad}(g)|_{\mathfrak{g_{u}}})=1\}$
So, $N$ normalizes $G_{u}$. Since $G$ is unimodular, by (5.15), $G\subseteq N$.
We have $\varphi:[0,\infty)\to G$. Let $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a correcting curve for $\varphi$, and denote $\psi(t):=\beta(t)\varphi(t)$. Now $N$ is observable in $\text{SL}(n,\mathbb{R})$ (Definition 2.2), and $\varphi([0,\infty))\subseteq G\subseteq N$. So $H_{\varphi}\subseteq N$ by the minimality of the hull, Theorem 2.5,(1). Now, by Remark 2.8, we have that $\psi([0,\infty))\subseteq H_{\psi}=H_{\varphi}$.
Assume for contradiction that $\psi$ is not bounded modulo $G_{u}$. Let $q:N\to N/G_{u}$ be the quotient map. Then, by Proposition 5.12, there exists a one-parameter unipotent subgroup of $\{u(s)\}_{s\in\mathbb{R}}\subseteq N$ such that
$q(u(s))=\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1},\,\forall s\in\mathbb{R},$
and $q\circ u$ is a nontrivial unipotent one-parameter subgroup of $N/G_{u}$. Let $\beta_{\infty}=\lim_{t\to\infty}\beta(t)$. Then, for all $s\in\mathbb{R}$,
@@
and since $\varphi([0,\infty))\subseteq G\subseteq N$, we get that $\{q(\beta_{\infty}^{-1}u(s)\beta_{\infty}):s\in\mathbb{R}\}$ is a nontrivial one-parameter subgroup of $G/G_{u}$. It then follows that $\{\beta_{\infty}^{-1}u(s)\beta_{\infty}:s\in\mathbb{R}\}$ is a nontrivial one-parameter unipotent subgroup contained in $G$ not contained in $G_{u}$, a contradiction. Finally, since $\varphi$ is bounded modulo $G_{u}$, and since $\text{zcl}(G_{u})$ is observable, we get that $H_{\varphi}\subseteq G_{u}$ by the minimality of the hull, Theorem 2.5,(1). ∎
6. Avoidance of singular sets
Suppose that $\Gamma\subseteq\text{SL}(n,\mathbb{R})$ is a discrete subgroup, and $\mu$ is a probability measure on $X:=\text{SL}(n,\mathbb{R})/\Gamma$ which is invariant under a nontrivial one-parameter unipotent subgroup of $\text{SL}(n,\mathbb{R})$. Then, by the celebrated Ratner’s theorem (see [Ratner91a]), every $W$-ergodic component of $\mu$ is homogeneous.
The measure $\mu$ that we will consider is a weak-* limit of the measures $\mu_{T,\varphi,x}$ (see (1.1)), for a non-contracting, definable polynomially bounded curve $\varphi$. By Theorem 4.5 and Proposition 5.1, $\mu$ is a probability measure, and it is invariant under a nontrivial unipotent subgroup.
As an immediate consequence of Ratner’s theorem (see Theorem 6.3) if $\mu$ is not $\text{SL}(n,\mathbb{R})$-invariant, then it is strictly positive on a Ratner’s singular set, which is the image of a certain strictly lower dimensional algebraic subvariety of $\text{SL}(n,\mathbb{R})$ on $X=\text{SL}(n,\mathbb{R})/\Gamma$. To analyze this situation, we will consider a finite-dimensional representation $V$ of $\text{SL}(n,\mathbb{R})$ and represent the singular set via an algebraic subvariety, say $\mathcal{A}$, in $V$. The trajectory $t\mapsto\varphi(t)x$ defining $\mu$ will be represented by countably many simultaneous trajectories $t\mapsto\varphi(t)\gamma\cdot p$ ($\forall\gamma\in\Gamma$) for a well chosen $p\in V$ with $\Gamma\cdot p$ discrete. We will show that $\mu$ cannot be positive on the singular set unless for some $\gamma\in\Gamma$, $\varphi(t)\gamma\cdot p$ converges to a point of $\mathcal{A}$ as $t\to\infty$. This leads to group-theoretic constraints.
Studying the dynamics of $\varphi(t)x$-trajectories in thin neighborhoods of a singular set in $X$ via simultaneously analyzing countably many trajectories of $\varphi$ in a thin neighborhood of an algebraic variety in a vector space is known as the linearization technique. Showing that an unbounded trajectory of $\varphi$ avoids remaining too close to an algebraic variety for too long is called the avoidance of singular sets - this is achieved via the $(C,\alpha)$-good property. The remainder of the section is devoted to employing these techniques in our situation.
6.1. Description of Ratner’s singular sets